OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

4x-12=24 4x=36 x=9

OpenStudy (anonymous):

check by plugging it in 8(9)-12=4(9)+24 60=60